Lock acquisition has gone bad tonight.
The initial stage works fine, up through handing off control of CARM to MCL. However, when increasing the AO path (analog gain), there are large DC shifts in the C1:IOO-MC_F signal. Eventually this causes the pockels cell in the FSS loop to saturate, and lock is lost.
This problem has disappeared. I don't know what it was.
The first plot shows one of the symptoms. The second plot is a similar section taken from a more normal acquisition sequence the night before.
All is not perfect, however, as now the handoff to RF CARM is not working.
The problem has returned. I still don't know what it is, but it's making me angry.